OpenStudy (anonymous):

The ratio of two numbers is 5 to 3. The sum of the numbers is 16. What are the numbers?

OpenStudy (anonymous):

\[\frac{ a }{ b }=\frac{ 5 }{ 3 },a+b=16\]

OpenStudy (anonymous):

Take some arbitrary number x. Then one number wud be 5x , while the other wud be 3x. Given that the sum of the numbers is 15, 5x + 3x =16 so 8x=16, which gives x=2 Now u can find the numbers ... 5x= 10 , 3x=6